OpenStudy (anonymous):

anita has two sisters and three brothers. the mean of all thier ages is 6 years. what will thiers mean age be 10 and 20 years from now

OpenStudy (anonymous):

15

OpenStudy (anonymous):

i think that's right but not sure

OpenStudy (anonymous):

thanks very much

OpenStudy (anonymous):

its not nvm

OpenStudy (anonymous):

sorry dont use it

OpenStudy (anonymous):

median increases by +10 i think so its 16 +20 i think so its 26 i think thats right

OpenStudy (anonymous):

yes Hahd looks right. S=sum of the 6 kid's ages and (S/6)=6 (the average is 6) Then 10 years later the sum of their ages will increase by 60, so the new average will be ((S+60)/6) = S/6+60/6 = 6+10 = 16 Same for 20 years later you get an average age of 26

OpenStudy (anonymous):

Since all the six people age simultaneously, their mean (average) age goes up with amount of years we go forward in time. This can be seen also from the expression of the average age. Whatever the ages \(a_1\), \(a_2\), \(a_3\), \(a_4\), \(a_5\) and \(a_6\) are, their calculated mean age is \(\dfrac{a_1+a_2+a_3+a_4+a_5+a_6}{6}\). If all ages get a +10 added to them, then the expression is \(\dfrac{(a_1+10)+(a_2+10)+(a_3+10)+(a_4+10)+(a_5+10)+(a_6+10)}{6}\), where the tens can be regrouped as follows: \[\dfrac{a_1+a_2+a_3+a_4+a_5+a_6+60}{6}\]\[=\dfrac{a_1+a_2+a_3+a_4+a_5+a_6}{6}+ \dfrac{60}{6}\]\[=\dfrac{a_1+a_2+a_3+a_4+a_5+a_6}{6}+ 10\]. Thus, the average age increased by the same amount of time as their individual ages.